1.- The probability that a computer is out of Fault service: type I is 1/5, type II is 13/50, type III is 23/100, of type I and II of 13/100, of type I and III of 3/20, of type II and III is 7/50 , and of all types 1/10, which Is the probability that it fails due to any of these causes? Answer:

2.- A person has the opportunity to win a prize, if he is able to strike with a soccer ball a circle of diameter 22 centimeters, equal to that of the ball. The circle is located in the center of a screen, placed in a goal whose dimensions are 560 centimeters by 200 centimeters. If any person can hit the space of the goal, what is the probability that a person will be able to Do you win the prize?

**please answer both. thankss
Answer #1

1 probability that it fails due to any of these causes =P(I or II or III)=P(I)+P(II)+P(III)-P(I and II)-P(I and III)-P(II and III)+P(I and II and III)

=(1/5)+(13/50)+(23/100)-(13/100)-(3/20)-(7/50)+(1/10)=37/100

2)

probability that a person will be able to win the prize =P(Area of center of screen)/P(total area)

=(pi/4)*222/(560*200)=0.003394